= 7. Determine whether the sequence an find the limit. (2n)3 +sin(n) n+n2 +6 converges or diverges. If it converges,
Answer #1

Claim: \lim_{n \to \infty }\frac{\sin n}{n^3}=0

We know,

0\leq \left | \frac{\sin n}{n^3} \right |= \frac{\left | \sin n \right |}{n^3}\leq\frac{1}{n^3}

\Rightarrow \lim_{n \to \infty }0\leq \lim_{n \to \infty }\left | \frac{\sin n}{n^3} \right | \leq \lim_{n \to \infty }\frac{1}{n^3}

\Rightarrow 0\leq \lim_{n \to \infty }\left | \frac{\sin n}{n^3} \right | \leq 0

sinn lim n+00 = 0 n3

\Rightarrow \lim_{n \to \infty } \frac{\sin n}{n^3} = 0

Given:

a_n=\frac{(2n)^3+\sin (n)}{n^3+n^2+6}

=\frac{8n^3+\sin (n)}{n^3+n^2+6}

  =\frac{8+\frac{\sin (n)}{n^3}}{1+\frac{n^2}{n^3}+\frac{6}{n^3}}

=\frac{8+\frac{\sin (n)}{n^3}}{1+\frac{1}{n}+\frac{6}{n^3}}

\Rightarrow \lim_{n \to \infty }a_n=\lim_{n \to \infty }\left ( \frac{8+\frac{\sin (n)}{n^3}}{1+\frac{1}{n}+\frac{6}{n^3}} \right )

  =\left ( \frac{8+\lim_{n \to \infty }\frac{\sin (n)}{n^3}}{1+\lim_{n \to \infty }\frac{1}{n}+\lim_{n \to \infty }\frac{6}{n^3}} \right )

=\left ( \frac{8+0}{1+0+0} \right )

  =8

Since, limit of the sequence \left \{ a_n \right \}^{\infty}_{n=1} exists hence, the sequence \left \{ a_n \right \}^{\infty}_{n=1} is convergent and its limit is 8.

Hope this helps!